OpenStudy (meehan98):

A die is rolled twice. What is the probability of getting either a multiple of 3 on the first roll or a total of 8 for both rolls?

OpenStudy (meehan98):

This is what I have so far: P(multiple of 3)=2/6 P (sum of 8)=6/16 Then, you add both of them to get 31/48, but this is incorrect.

OpenStudy (anonymous):

Wait I am not sure. Lemme think about it more

OpenStudy (anonymous):

What is the correct answer?

OpenStudy (meehan98):

Well, these are the possible answers: 5/12 7/18 13/180 4/9

OpenStudy (meehan98):

Yea, I got 25/108 so something's off; I just don't know what?

OpenStudy (anonymous):

Oh, I read the question wrong. I thought it was probability of of getting multiple of 3 and getting a roll of 8 on both rolls.

OpenStudy (meehan98):

So, I keep getting different answers each time I complete the problem. Now I got 8/9 because I added 2/6 + 5/6=7/6 and subtracted that from 10/36 according to the Inclusive events formula P ( A or B)= P(A) +P(B)-P(A and B)

OpenStudy (anonymous):

Its been a while. Gimme a few more minutes.

OpenStudy (meehan98):

No, you're fine!

OpenStudy (anonymous):

Got it. P(A)+P(B)- P( A and B)

OpenStudy (anonymous):

For the multiple of 3 only three and six work.

OpenStudy (anonymous):

2,6 3,5 4,4 5,3 6,2 All this would add up to 8. Now we would use independent event formula by multiplying probability by probability The chance of getting any of those number is

OpenStudy (anonymous):

Equation bugged out.

OpenStudy (anonymous):

Continued from 2 post above is \[\frac{ 5 }{ 6 }\] Chances for second dice would be 1/6 since you only need one number to make it 8. So \[\frac{ 5 }{ 6 } \times \frac{ 1 }{ 6 }=\frac{ 25 }{ }\]

OpenStudy (anonymous):

5/36*

OpenStudy (anonymous):

(5/36+2/6) -P(A and B) P(A and B) would be \[\frac{ 2 }{ 6 } \times \frac{ 1 }{ 6 }=2/36\]

OpenStudy (anonymous):

You could have wait. If it was a timed test, you should have studied -_-

OpenStudy (meehan98):

I understand how you have received all of these besides one part. (5/36+2/6) - P (A and B) I know that it's 2/6 from the multiple of 3 but where did you get 1/6? because I thought that P(total of 8 for both rolls)=5/36.

OpenStudy (meehan98):

It wasn't a test, it's a lesson that I didn't fully grasp. I apologize for not responding sooner; I was working out the problem again. You've been wonderful help. Thank you!

OpenStudy (anonymous):

Ah, I see. I was eating dinner, so sorry for the delay. P(A+B) is the probability for both event (A&B) to happen. A is the probability for it to be multiples of 3 which is 2/6 B is probability to add up to 8. If you rolled 3 on first dice, then second dice would need to be 5 to make it 8. If you rolled 6 on first dice, second dice would need to be 2 to make it 8. Because even if you were lucky to roll a 3 on first dice, you would need 5 on second dice. That would be a 1/6 chance so 2/6 x 1/6

OpenStudy (meehan98):

Ah, you're wonderful at explaining things! Thank you so much!!!

OpenStudy (anonymous):

Thanks for the feedback and nice job attempting it. Good luck on the rest!